# LoQT: Low Rank Adapters for Quantized Training

### v1

@@ 4 Experiments and Results
4.1 Experimental Setup
We evaluate LoQT by training LLaMA-based[34] language models on the C4 dataset [30], a collection of processed text in English that was scraped from the internet [30]. We train models of sizes of 60M, 130M, 350M, and 1B parameters, adhering to single-epoch training cycles determined by Chinchilla Scaling Laws [15]. While LoQT is capable of training models up to 13 billion parameters on consumer GPUs, compute limits prevent us from training to convergence for sizes above 1B. We also benchmark LoQT on the GLUE test-suite for natural language understanding [37]. Runs were conducted on up to 4x 40GB NVIDIA A100s 2x 80GB NVIDIA H100s, or a single 24GB NVIDIA RTX 3090. The longest run was the training of the 1B models, taking approximately four days on the four A100s. The 3090 was used for throughput and to empirically verify memory claims.
Hyperparameters are consistent across model sizes, with experiments in BF16 format for memory efficiency. All models use a maximum sequence length of $256$, a total token batch size of 131K tokens, and a learning rate warmup for the first 10% of the training steps, followed by cosine annealing to 10% of the initial learning rate. Full experimental details, including the specific hyperparameters for each task, are provided in Appendix B.
@@ 4.1 Experimental Setup
For pre-training, we compare LoQT against LoRA [16], ReLoRA [22], GaLore [44], and non-quantized version of LoQT, LoQT-nq. In our experiments, LoQT, LoRA, and ReLoRA modify attention and fully connected layers while maintaining full-rank embeddings and normalization layers. This contrasts with GaLore, which keeps weights full-rank but projects gradients to low-rank, and standard full-rank training. For fine-tuning, we benchmark LoQT against GaLore, LoftQ [21], LoRA and LoQT-nq. All models use identical update frequencies for GaLore, ReLoRA, LoQT-nq, and LoQT, starting with an update frequency of $T=100$ and then with exponentially increasing update frequencies. This means that we do more frequent updates early and fewer as the model stabilizes (see Section 4(b) for more details). All models are trained using the Adam optimizer, except GaLore which uses GaLoreAdam for gradient projection.
4.2 Pre-training of Generative Language Models
Results and details for pretraining of language models of sizes 60M, 130M, 350M and 1B parameters are shown in Table 1. Model sizes are calculated based on the full models without any low-rank methods. We see that LoQT and LoQT-nq both perform very close to full rank pretraining and GaLore, while using significantly less memory by keeping most of the model weights in a quantized state. For the 60M model, full training is only slightly better than LoQT, while we see results improve or be within the standard error for the other sizes. We also notice a slight drop in performance from quantizing the original weight matrix, comparing LoQT and LoQT-nq. The key difference between the approaches are the theoretical memory estimaes, e.g. where LoQT requires 59% less memory for the 1B model in full precision and 28% less memory than GaLore.
Table 2: Results with LoQT, LoQT-nq, and GaLore of DeBERTaV3-base models on the GLUE development set. We report mean and standard error over three seeds. The best results on each dataset are shown in bold.
@@ 4.3 Memory-efficient finetuning
We fine-tune the pre-trained DeBERTa-V3-base [^1] From https://huggingface.co/microsoft/deberta-v3-base. [13] model on GLUE tasks using LoQT and compare its performance with a full fine-tuning baseline, LoRA, LoftQ, and GaLore. See Appendix 5 for details on hyperparameters. Results are given in Table 2.
We find that both LoQT-nq and LoQT perform well. And somewhat surprisingly, it sometimes surpasses GaLore, LoftQ, and LoRA.This may indicate that initializing the LoRA factors with information about the gradient of $W$ could be a beneficial starting point compared to standard initialization methods. As the goal of this work is to limit memory consumption, we leave out further comparisons that could verify these findings to future work.
@@ 4.4 Memory and Throughput
Table 3: Comparison of memory usage between GaLore, LoRA, and LoQT. $W\in\mathbb{R}^{m\times n}$ ($m\leq n$), rank $r$.
@@
An overview of memory usage for GaLore, LoRA and LoQT is given in Table 3. We see that LoQT makes use of the same number of trainable parameters as LoRA for a given rank while using less memory for the optimizer states and gradients than in both LoRA and GaLore.
We compare the LoQT to the closest in memory performance approach, GaLore, for 13B in Figure 1, and for other model-sizes in Figure 6. We compare three different use cases, using the approaches directly, combining them with an 8-bit Adam optimizer [5], and using per-layer weight updates with offloading (while still using 8-bit Adam). We see from the figures that LoQT significantly shrinks both the number of trainable parameters and optimizer states compared to GaLore.
Per-layer weight update is essential for GaLore; without it, an additional $\sim\!\!12$ GB of VRAM is needed for gradients in a 7B model, making full-parameter fine-tuning impossible on a 24GB GPU. Additionally, the per-layer gradient updates may not work well with DDP (Distributed Data Parallel) and gradient accumulation. With our method, we can get a lower memory than GaLore even when they use per-layer gradient updates. When not using per-layer gradient updates, this difference becomes even bigger as seen for the 7B model in Figure 6.
Moreover, our method supports training 7B models without per-layer computations on 24GB GPU. This makes it possible to use multi-GPU training, a capability not possible with the current GaLore approach. Our memory advantage allows for a batch size of $1280$ tokens compared to GaLore’s $256$ for the 7B model on the 24GB RTX3090. With per-layer gradient updates, LoQT can train a 13B model on a single GPU, pushing the limits of hardware efficiency.
@@
We evaluate the throughput with a sample batch size of 16 with a total batch size of 512 using gradient accumulation, which is the largest power of 2 that fits on the GPU. We update the projection matrix $P$ for every 200 iterations. The per-layer gradient update algorithms apply a weight update for every mini-batch as they do not support gradient accumulation. For the evaluation, we use a 1B parameter model with rank 512. We find that LoQT can process 16% fewer tokens per second than only using AdamW, at 3996 tokens/s compared to 4782 tokens/s on the RTX3090.
@@ 5 Ablations
(a) EC: Error compensation, EF: Exp. decreasing update frequency.
(b) Ablation for different update frequencies, with exponentially increasing at the bottom
Figure 4: Ablation results for update frequency, error-compensation, quantization, model size 130m, and rank $256$.
@@
To assess the impact of quantization error compensation, we analyze the validation loss curves for a 130 million parameter model. Figure 4(a) shows that quantizing $W$ or both $W$ and $P$ without error compensation, or exponential frequency updates, causes the loss to stagnate early. We also note that quantizing $P$ has a much smaller effect on the loss compared to quantizing $W$. Error compensation significantly improves the model’s performance, resulting in approximately 3.5 points better perplexity. Adding exponentially increasing update frequency improves perplexity by an additional 1.5 points, achieving performance close to that of models without quantization.
@@
Without the quantization error compensation detailed in §3.5, LoQT’s performance stagnates earlier and diverges more from the other models. This demonstrates the effectiveness of our compensation approach in mitigating the quantization errors introduced during the $W$ update with $AB$ and subsequent quantization steps.
Projection update frequency
@@
Our scheduling approach ensures more frequent updates early in training to facilitate substantial weight adjustments. As training progresses, the update frequency decreases, allowing for the accumulation of larger updates to compensate for smaller updates that might be canceled out by quantization errors. Figure 4(b) presents an ablation study on our method of progressively increasing update frequency starting at $100$ and increasing by a factor of $1.2^{T}$ up to $2500$. We show the validation loss curves for fixed update frequencies $200$, $400$, $500$, and $1000$.
@@
The results show that exponentially increasing the update gap is particularly beneficial for models employing quantization, enabling them to achieve the same perplexity as those without quantization while making use of GaLore. Conversely, the performance gains are more subtle for models that do not use quantization and rely solely on GaLore. We hypothesize that even these models might benefit from the larger projection gap intervals. This could be due to the reduction in the accumulation of errors from frequent updates of the projection factor $P$, as the influence of outdated optimizer statistics becomes less prevalent. Finally, an ablation on the ranks used for $P$ and $B$ is given in Figure 5 in the Appendix.
@@ 6 Discussion and Conclusion
We present LoQT, a method for memory-efficient pretraining and adaptation of quantized models. The key insights behind the approach are the benefits of initializing low-rank factors using the gradient of the weight matrix and using exponentially increasing update gaps that make updating of a quantized model possible. While our initial goal was to lower memory usage to facilitate the training of models such as LLMs on consumer-grade hardware, we are cautiously excited about the results sometimes being better than the baselines. These evaluations will be explored in more detail in future work.
Our method is general and opens up new ways of decreasing memory use as well as improving the training throughput. This could be done by implementing kernel fusion and using other quantization methods such as NF2 [6] or quantization of activations, making it possible to do the matrix multiplications using modern tensor core formats such as FP8 or INT4.
@@ 7 Impact and Limitations
Our work has the potential to have a significant impact on those working in hardware-constrained settings by enabling more efficient training on consumer hardware. We are particularly excited to see the method being applied in single GPU settings. We validate LoQT on several model sizes, by training over many steps and by fine-tuning on a standard benchmark for natural language understanding. While we are confident in our results, further exploration of training duration, data diversity, and hyper-parameter tuning might lead to different results in those settings.
@@ 8 Acknowledgements
This work is supported by the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516). Serge Belongie and Vésteinn Snæbjarnarson are supported by the Pioneer Centre for AI, DNRF grant number P1. MJK acknowledges support from the Carlsberg Foundation and the Novo Nordisk Foundation.
@@ Appendix A Quantization Methods
PQT [9, 7, 35, 42, 27, 4, 20, 14, 31] involves converting a pre-trained high-precision model into a lower precision format. This can be done directly or by using a subset of the training data to calibrate the quantization process or fine-tune the quantized weights to adapt the model to the quantization. However, PTQ often results in reduced accuracy compared to QAT because the model does not learn to adjust to quantized weights during training [9, 42].
@@ Appendix B Hyperparamters
B.1 Pre-training
For the pre-training results shown in Table 1, we adopted configurations from GaLore [44] and tested pre-training methods on different LLaMA 2 model sizes using the C4 dataset. Training was conducted with optimizer states in BF16 precision, and NF4 precision quantization was used for LoQT. The model rank was adapted based on the largest layer with specific parameters.
Table 1 shows the ratio $r/d_{model}$, which denotes the rank relative to the largest weight matrix dimension. All experiments used a maximum sequence length of 256, learning rate warmup for the first 10% of training steps, and cosine annealing for the learning rate schedule, decaying to 10% of the initial rate.
Galore, LoQT-nq, and LoQT used exponentially increasing update frequencies starting at 100 and increasing by $100+\psi^{i}$, where $\psi$ is 1.2 and $i$ is the update counter (see Section C.1 for more details).
We tested learning rates of $0.01$, $0.005$, $0.001$, and $0.0005$ across different model sizes. For models ranging from 60M to 350M parameters, a learning rate of $0.01$ yielded the best performance. In contrast, full-rank models required smaller learning rates: $0.001$ for 60M to 350M models and $0.0005$ for the 1B model. To scale the learning rates for LoQT, LoQT-nq, and Galore, we employed a scale parameter $s$ set to $0.5$ and $0.25$ for the 1B model. This parameter functions similarly to the LoRA alpha parameter, determining the weight on the learned factors for LoQT and LoQT-nq. For Galore, our experiments indicated that $s=0.5$ was more effective than the $0.25$ reported in [44]. This scaling approach effectively adjusts the learning rate, resulting in an actual rate of $0.005$ for the multi-head attention and feed-forward layers in LLaMA models, which is relatively large compared to the $0.001$ used for full-rank models. Higher learning rates led to spikes in the training loss for both full-rank and LoQT models.
Table 4: Pre-training hyperparameters of LLaMA models for evaluation. (-) Indicates we did not train such a model.

### v4

@@ 3 Experiments
We evaluate LoQT on language model pretraining by training LLaMA-based [15] language models on the C4 dataset [16], a collection of text in English that was extracted from the Common Crawl web-scrapes [16]. We train models of sizes of 60M, 130M, 350M, and 1B parameters, adhering to single-epoch training cycles determined by the Chinchilla Scaling Laws [17]. While LoQT is capable of training models up to 13 billion parameters on consumer GPUs, compute limits prevent us from training to convergence for sizes above 1B. We also benchmark LoQT on the GLUE test-suite for natural language understanding [18], the GSM8K [12] dataset for arithmetic reasoning and an Icelandic text dataset [10] to evaluate language adaptation via continued-pretraining. Runs were conducted on up to 4x 40GB NVIDIA A100s 2x 80GB NVIDIA H100s, or a single 24GB NVIDIA RTX 3090. The longest run was the training of the 1B models, taking approximately four days on the four A100s. The RTX 3090 was used for throughput and to empirically verify memory claims.
We keep hyperparameters consistent across model sizes, with experiments conducted in BF16 format for memory efficiency. All models are trained with a maximum sequence length of $256$, a total token batch size of 131K tokens, and a learning rate warmup for the first 10% of the training steps, followed by cosine annealing to 10% of the initial learning rate. Full experimental details, including the specific hyperparameters for each task, are provided in Appendix B.
@@
For pretraining, we compare LoQT against LoRA [5], ReLoRA [7], GaLore [8], and a non-quantized version of LoQT, LoQT-nq. In our experiments, we apply these parameter-efficient training methods to the attention projection matrices and fully connected layers while maintaining full-rank embeddings. For the fine-tuning experiments, we compare LoQT against GaLore, LoftQ [14], LoRA, ApiQ [4], and LoQT-nq, or a subset thereof. All models that make use of update frequencies are trained using the same intervals, these are GaLore, ReLoRA, LoQT-nq, and LoQT. We start with an update interval of $T=100$ and then exponentially increase the update frequency. This means that we do more frequent updates early and fewer as the model stabilizes (see § 4(b) for more details). A scaling parameter $\alpha=0.5$ is used for LoQT and GaLore across all models, except for the 1B model where it is decreased to $0.25$. The same rank $r$ is used for all low-rank methods. All models are trained using the Adam optimizer, except GaLore which uses their GaLoreAdam optimizer for gradient projection. More details on hyperparameters are provided in the Appendix B.
3.1 Pretraining of Generative Language Models
Results and details of pretraining causal language models of sizes 60M, 130M, 350M, and 1B parameters are shown in Tab. 1. Model sizes are calculated based on the full models without any low-rank methods. We see that LoQT and LoQT-nq both perform very close to full rank pretraining and GaLore while using significantly less memory by keeping most of the model weights in a quantized state. For the 60M model, full training is only slightly better than LoQT, while we see results improve or stay within the standard error for the other sizes. We also notice a slight drop in performance from quantizing the original weight matrix, comparing LoQT and LoQT-nq. The key difference between the approaches is the theoretical memory estimates, e.g. where LoQT uses 59% less memory for the 1B model in full precision and 28% less memory than with GaLore.
Table 2: Results for LoQT, LoQT-nq, and GaLore using DeBERTaV3-base models on the GLUE development set. We report mean and standard error over three seeds. The best mean results on each dataset are shown in bold.
@@ 3.2 Memory-Efficient Finetuning
We fine-tune the pretrained DeBERTa-V3-base [^1] From https://huggingface.co/microsoft/deberta-v3-base. [19] model on the natural language understanding GLUE [11] tasks using LoQT and compare its performance with full fine-tuning baselines, LoRA, LoftQ, and GaLore. See Appendix 7 for details on hyperparameters. Results are given in Tab. 2.
We find that both LoQT-nq and LoQT perform well. And somewhat surprisingly, they sometimes surpass GaLore, LoftQ, and LoRA. This may indicate that initializing the LoRA factors with information about the gradient of $W$ is a beneficial starting point compared to standard initialization methods. Further experiments are needed to confirm and investigate these findings which we leave to future work.
Arithmetic Reasoning on GSM8K
We fine-tune quantized Llama-2 models (7B and 13B) on the GSM8K dataset [12] for arithmetic reasoning. As shown in Tab. 4, LoQT achieves average test set accuracies of $42.6\%$ and $52.9\%$ with the 7B and 13B models, respectively, performing comparably to other quantized fine-tuning approaches. Detailed hyper-parameters are provided in Appendix Tab. 8.
Table 3: GSM8K LLaMA-2 7B and 13B test accuracy with std. error. Best mean is in bold.
Method Bit LLaMA-2-7B LLaMA-2-13B
LoRA 16 41.7 $\pm$ 0.3 51.3 $\pm$ 0.86
QLoRA 4 41.9 $\pm$ 0.2 51.6 $\pm$ 0.29
LoftQ 4 41.9 $\pm$ 0.9 51.3 $\pm$ 0.96
ApiQ 4 42.1 $\pm$ 0.5 52.4 $\pm$ 0.46
LoQT 4 42.6 $\pm$ 0.4 52.9$\pm$ 0.12
Table 4: Llama-7B fine-tuning on Icelandic. We report test set perplexity.
Method Perplexity $\downarrow$
No training 4.90
Full 3.79
GaLore 3.96
LoQT-nq 3.61
LoQT 3.63
Continued Pretraining of Llama 7B
We also evaluate LoQT on language adaptation of a large language model. We continue pretraining of the Llama-2-7B model using a curated subset of a public Icelandic text dataset extracted from [10] containing  770k documents. We compare LoQT with NF4 quantization, LoQT without quantization (LoQT-nq), regular training, and GaLore, using consistent hyper-parameters across all methods, results are shown in Tab. 4. LoQT achieves test set perplexity close to that of using full training or GaLore while reducing perplexity from 4.90 (non-trained model) to 3.63. Additional details are provided in Appendix C.1.
@@ 3.3 Memory and Throughput
Table 5: Comparison of memory usage for GaLore, LoRA, and LoQT. $W\in\mathbb{R}^{m\times n}$ ($m\leq n$), rank $r$.
@@
An overview of memory usage for GaLore, LoRA and LoQT is given in Tab. 5. We see that LoQT has the same number of parameters as LoRA for a given rank while using less memory for the optimizer states and gradients than in both LoRA and GaLore.
We compare LoQT to GaLore, the approach that gets closest in memory performance, for a model of size 13B in Fig. 1, and for other model-sizes in Fig. 6. We compare three different use cases, applying the methods on their own, combining them with an 8-bit Adam optimizer [20], and using per-layer weight updates with offloading (while still using 8-bit Adam). We see that LoQT significantly reduces the number of trainable parameters, and optimizer states, compared to GaLore.
Per-layer weight updates are essential for GaLore; without it, an additional $\sim\!\!12$ GB of VRAM is needed for the gradients of a 7B model, making full-parameter fine-tuning impossible on a 24GB GPU. Additionally, the per-layer gradient updates do not work with gradient accumulation. Using LoQT results in lower memory use than GaLore, even with per-layer gradient updates. When not using per-layer gradient updates, the difference becomes more pronounced as seen for the 7B model in Fig. 6.
LoQT enables training of 7B models without per-layer computations on a 24GB GPU, allowing for gradient accumulation and higher effective batch sizes. Our memory advantage allows for a batch size of $1280$ tokens compared to GaLore’s $256$ for the 7B model on the 24GB RTX 3090. Using per-layer gradient updates, LoQT can train a 13B model on a single GPU. We refer to Fig. 8 in the Appendix for a comparison of how Adam, GaLore, and LoQT scale with increasing context length.
@@
We evaluate the throughput with a sample batch size of 16, and a total batch size of 512 using gradient accumulation, which is the largest power of two that fits on the GPU. We update the projection matrix $P$ for every 200 iterations. We evaluate the throughput using a 1B parameter model and rank 512 without per-layer gradient updates. We find that LoQT processes 16% fewer tokens per second than training with only AdamW, at 3996 tokens/s compared to 4782 tokens/s on the RTX 3090.
@@ 4 Ablations
(a) EC: Error compensation, EF: Exp. increasing update interval.
(b) Ablation of update intervals: comparing fixed intervals to an exponentially increasing schedule.
Figure 4: Ablation results for update intervals, error-compensation, quantization using model size 130m, and rank $256$. $W_{q}$: quantized $W$; $P_{q}$: quantized $P$; No Q: no quantization. The dynamic update interval $100+1.2^{i}$ grows exponentially for each step $i\in\mathbb{N}$.
@@
We analyze the validation loss curves of $130$ million parameter models to assess the impact of error quantization compensation. Fig. 4(a) shows that quantizing $W$, or both $W$ and $P$, without error compensation, or exponential interval updates leads to early stagnation of the loss. We also note that quantizing $P$ has a much smaller effect on the loss compared to quantizing $W$. Error compensation significantly improves the model’s performance, resulting in approximately $3.5$ points better perplexity. Adding exponentially increasing update intervals improves perplexity further by an additional $1.5$ points, achieving performance close to that of models without quantization.
@@
Without the quantization error compensation, detailed in §2.5, LoQT’s performance stagnates earlier and diverges more from the other models. This demonstrates the effectiveness of our compensation approach in mitigating the quantization errors introduced during the update of $W$ with $PB$ and subsequent quantization steps.
Projection Update Intervals
@@
Our scheduling approach ensures more frequent updates earlier on in training when the weight adjustments are larger. As training progresses, the update intervals get larger, allowing for accumulating more updates to compensate for smaller changes at each step that might otherwise be canceled out by the quantization errors. Fig. 4(b) presents an ablation study of progressively increasing update intervals starting at $100$ and increasing by a factor of $1.2^{T}$ up to $2500$. We show the validation loss curves for fixed update frequencies $200$, $400$, $500$, and $1000$.
@@
The results show that exponentially increasing the update interval is particularly beneficial for models employing quantization, enabling them to achieve the same perplexity as those without quantization. Conversely, the performance gains are more subtle for models that do not use quantization. We hypothesize that even these models might benefit from the larger projection interval intervals. This could be due to the reduction in the accumulation of errors from frequent updates of the projection factor $P$, as the influence of outdated optimizer statistics becomes less prevalent. Finally, an ablation on the ranks used for $P$ and $B$ is given in Fig. 5 in the Appendix.
5 Related Work
We now provide an overview of related work on quantization, parameter-efficient fine-tuning methods, and memory-efficient approaches.
5.1 Neural Network Quantization and NF4
$w_{\text{NF4}}$ $=q_{\text{NF4}}(w,\mathcal{M}_{\beta})$ (9)
[…]
$w$ $=q_{\text{NF4}}^{-1}(w_{\text{NF4}},\mathcal{M}_{\beta})$ (11)
[…]
5.2 Adaptation of Pretrained Networks
Low-Rank Adaptation (LoRA) [5] enables fine-tuning of pretrained models using low-rank adaptors, effectively reducing the memory footprint by only training weight adaptors for targeted layers. However, simple low-rank training using LoRA factor matrices has not been shown to work for pretraining [7].
5.3 Memory Efficient Optimization
Optimizer memory consumption
A significant portion of the memory needed to train neural networks is typically consumed by optimizer states. Notably, Adam [9], one of the most widely used optimizers, uses double the amount of memory as the gradient matrix to maintain first and second-order gradient statistics. Efforts to reduce this overhead have led to the development of adaptive optimization algorithms like Adafactor [24], which achieves sub-linear memory costs by factorizing the second-order statistics into a row-column outer product. GaLore [8] expands on this concept by using low-rank factorization and projecting low-rank gradients up to full size when updating model weights.
Periodic updating of weight matrices
Low-rank Gradients
@@ 6 Discussion and Conclusion
We have presented LoQT, a method for memory-efficient pretraining and adaptation of quantized models. Key insights behind the approach are the benefits of initializing low-rank factors using the gradient of the weight matrix and using exponentially increasing update gaps that make updating of a quantized model feasible. While our initial goal was to lower memory usage, to facilitate the training of models such as LLMs on consumer-grade hardware, we are also cautiously excited about the results sometimes exceeding those of the baselines. We hope to see this explored in more detail in future work.
Our method is general and has the potential to open up new ways of decreasing memory use and improving training throughput through further optimization of our implementation. This could be done by using other quantization methods such as NF2 [3] or quantization of activations, making it possible to do the matrix multiplications using modern tensor core formats such as FP8 or INT4.
@@ 7 Impact and Limitations
The presented work has the potential to benefit those working in hardware-constrained settings by enabling more efficient training on consumer-grade hardware. We are particularly excited to see the method being applied in single GPU settings.
We have validated LoQT on several model sizes, by training over many steps, by fine-tuning on a standard benchmark for natural language understanding, mathematical reasoning, and language adaptation. While we are confident in our results, further exploration of training duration, data diversity, and hyper-parameter tuning might lead to different results in those settings and we encourage users to confirm the benefit of LoQT for their approach.
@@ 8 Acknowledgements
This work is supported by the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516). Serge Belongie and Vésteinn Snæbjarnarson are supported by the Pioneer Centre for AI, DNRF grant number P1. MJK acknowledges support from the Carlsberg Foundation and the Novo Nordisk Foundation. Mads Toftrup gratefully acknowledges the Data-Intensive Systems research group at Aarhus University for providing GPU access.
@@ Appendix A Quantization Methods
PQT [31, 32, 33, 34, 35, 20, 36, 37, 38] involves converting a pretrained high-precision model into a lower precision format. This can be done directly or by using a subset of the training data to calibrate the quantization process or fine-tune the quantized weights to adapt the model to the quantization. However, PTQ often results in reduced accuracy compared to QAT because the model does not learn to adjust to quantized weights during training [31, 34].
@@
Choice of Quantization Method
We chose NF4-quantization because it has been shown to work well [46, 3]. Unlike other methods that adapt low-rank factors to quantization, such as LoftQ [14] and LQLoRA [47], LoQT operates under different constraints. Specifically, we do not have the flexibility to freely choose both $A$ and $B$ factors because the matrix $A$ is already fixed as the projection matrix $P$ containing gradient information. Both LoftQ [14] and LQLoRA [47] use the SVD of the quantization error to initialize the low-rank factors $A$ and $B$, aiming to minimize the difference between the quantized $W$ and $Q+AB$. The SVD gives the factorization $U\Sigma V^{T}$, where the top $r$ singular vectors in $U$ and $V$ are used to initialize the low-rank factors $A$ and $B$, respectively.
In contrast, LoQT takes a different approach due to the fixed nature of our low-rank adapter $P$ (analogous to $A$ in LoftQ and LQLoRA). Instead of applying SVD to the quantization error, we aim to minimize an objective where $W$, $Q$, and $P$ are fixed. We derive $B$ using the formula $B=P^{+}(W-Q)$, where $P^{+}$ is the Moore-Penrose pseudo-inverse of $P$ rather than the inverse. Incorporating information about the diagonal approximation of the Fisher information matrix into our objective could potentially reduce the error even further, a direction we are interested in exploring in future work.
@@ Appendix B Hyperparamters
We provide the hyperparameter configurations and setups used in our experiments to facilitate the reproduction of all results presented in this paper.
B.1 Pretraining
For the pretraining results shown in Tab. 1, we adopted configurations from GaLore [8] and tested pretraining methods on different LLaMA 2 model sizes using the C4 dataset. Training was conducted with optimizer states in BF16 precision, and NF4 precision quantization was used for LoQT. The model rank was adapted based on the largest layer with specific parameters.
Tab. 1 shows the ratio $r/d_{model}$, which denotes the rank relative to the largest weight matrix dimension. All experiments used a maximum sequence length of 256, learning rate warmup for the first 10% of training steps, and cosine annealing for the learning rate schedule, decaying to 10% of the initial rate. Galore, LoQT-nq, and LoQT used exponentially increasing update frequencies starting at 100 and increasing by $100+\psi^{i}$, where $\psi$ is 1.2 and $i$ is the update counter (see Section C.1 for more details).
We tested learning rates of $0.01$, $0.005$, $0.001$, and $0.0005$ across different model sizes. For models ranging from 60M to 350M parameters, a learning rate of $0.01$ yielded the best performance. In contrast, full-rank models required smaller learning rates: $0.001$ for 60M to 350M models and $0.0005$ for the 1B model. To scale the learning rates for LoQT, LoQT-nq, and Galore, we employed a scale parameter $s$ set to $0.5$ and $0.25$ for the 1B model. This parameter functions similarly to the LoRA alpha parameter, determining the weight of the learned factors for LoQT and LoQT-nq. For Galore, our experiments indicated that $s=0.5$ was more effective than the $0.25$ reported in [8]. This scaling approach effectively adjusts the learning rate, resulting in an actual rate of $0.005$ for the multi-head attention and feed-forward layers in LLaMA models, which is relatively large compared to the $0.001$ used for full-rank models. Higher learning rates led to spikes in the training loss for both full-rank and LoQT models.
Table 6: Pretraining hyperparameters of LLaMA models for evaluation. (-) Indicates we did not train such a model.